Tohban Report 2015-01-07

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 31 Dec 2014
End Date: 7 Jan 2015
Tohban: Milo Buitrago-Casas
Tohban email:
Next Tohban: Lindsay Erin Glesener
List all reports


Solar Activity

Solar activity was moderate this week with two M-class flares, several C-class flares and a handful of micro-flares that specially occurred at midlatitudes on Active Region 12253. This AR is still on the disk and has a good chance for a >M class flare. Currently there are 6 active regions on the disk, one is rotating off and two ARs will show up later today on the east limb.

How many GOES flares occurred?

 Flares above B, C, M, X class were      4    37     2     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      2    26     2     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 408 / 43 over the time range 30-Dec-14 06-Jan-15

Memory Management

The SSR was emptying by the end of the daily pass set during most of the week. Decimation was changed from Active/Vigorous to Normal/Vigorous on Dec 31, 2014 at 20:17 UTC. In addition, we changed night time data collection for the front segment to record data during whole eclipses. Thus, now RHESSI's front and rear segment data will remain on through eclipse starting on January 6, 2015 at 18:48 UTC.

Because this morning we changed G7 thresholds on the front fast and front slow and front fast value is showing very high number on monitor rates (around 900), we recommend to the next Tohban working with Albert to fix this issue. This fact could make that our SSR fills a little bit faster.

Spacecraft Status

Detectors 1,3,6, 8, and 9 are stably segmented. Detectors 5 and 7 are roughly segmented. This segmentation is expected to become more stable over time. Detectors 2 and 4 are still unsegmented.

We still see light leak on detectors 2, 4 and 9.

Data Gaps

There was no significant data gaps.


Cold plate temperatures are at 131.196 K and 129.341 K.

Detector issues

On January 5 at 14:57 UTC we changed thresholds on the front fast and front slow values for detector five [G5].

Threshold [G5] Front Fast Front Slow
Initial 0x60 0x80
Final 0x2E (same as initial G7) 0x10 (nominal value)

During the changes we tried to set the front fast threshold to the nominal value (0x22), but monitor rate in that case showed a very high value (around 700). Then we decided to set the front fast threshold to the same value that G7 initially had. Threshold front slow value was well-set to nominal. After these changes, monitor rates on G5 looked stable and in accordance with the other well-behaved segmented detectors (i.e. 1,3,6,8, and 9). Next Tohban should specially keep an eye on the behaviour of this detector.

D7 front fast and front slow were changed today January 7, 2015 at 14:09 UTC.

Threshold [G7] Front Fast Front Slow
Initial 0x2E 0x10
Final 0x22 (nominal value) 0x0C (nominal value)

Only detectors 2 and 4 continue unsegmented.

Light leak in D2 front, D9 rear and D5 rear.

Other notes

Next Tohban should keep an eye on thresholds for detectors 5 and 7, and regard changing threshold values for detector 9.

Spacecraft Management

Decimation Normal/Vigorous
Night time data (fronts) full nighttime events
Night time data (rears) full nighttime events
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools